OpenStudy (seattle12345):

Two servings of chicken and eight servings of turkey provide 298 grams of protein. Three servings of chicken and two servings of turkey provide 112 grams of protein. Determine the protein content in each item.

OpenStudy (lgbasallote):

let x = chicken y = turkey 2x + 8y = 298 3x + 2y = 112 solve..you know how?

OpenStudy (seattle12345):

2x + 8y = 298 3x + 2y = 112 2x=-8y+298 X=-4y+149 3(-4y+149)+2y=112 -12y+447+2y=112 -10=-335 Y=35.5 ?

OpenStudy (anonymous):

its right.y=33.5

OpenStudy (lgbasallote):

335/10 = 3.35 the rest..you're clean ^_^

OpenStudy (anonymous):

?????????????

OpenStudy (seattle12345):

wouldn't it be 33.5?

OpenStudy (lgbasallote):

yeah 33.5...typo lol sorry

OpenStudy (seattle12345):

haha its cool. thanks for your help guys!!!

OpenStudy (lgbasallote):

i think your'e supposed to solve for x too

OpenStudy (seattle12345):

oh yea, I did. x=15

OpenStudy (lgbasallote):

sounds right